# Hovering Flight in Flapping Insects and Hummingbirds: A Natural Real-Time and Stable Extremum Seeking Feedback System

### v3

@@ III.1 Results and Simulation
In equilibrium hovering, the velocity $w$ has to be 0 as this translates into constant height/altitude $z$ (see the first differential equation in (11)). Therefore, in our simulations we start with an initial disturbance in $w$, that is $w_{0}=0.2$. This disturbance challenges the natural hovering ES system as it needs to overcome temporary deviations from equilibrium and re-establish hovering equilibrium (i.e., $w=0$), demonstrating stability in real-time. The simulations were performed using two objective functions: $J=z^{2}$, where its minimum represents the hovering condition (constant $z$) and $J=\dot{w}^{2}$, where its minimum represents the hovering condition (lift to weight ratio around one). By employing both objective functions in our simulations, the system demonstrated an impressive ability to re-stabilize itself, with the vertical velocity state $w$ oscillating around zero and the altitude remaining constant over time. This behavior reflects stable return to equilibrium hovering after disturbance. The state variables response of the natural hovering ES system (11) are provided in Figure 5 for the hawkmoth, and Figures 12, 13, 14, 15 and 16 in the Appendix for the other insects and the hummingbird. The left column of each figure corresponds to results under the objective function $J=z^{2}$, while the right column of each figure corresponds to results under the objective function $J=\dot{w}^{2}$. The stable hovering conditions of our simulation results (constant $z$ and lift to weight ratio of one) are further clarified in Figure 6 for hawkmoth and Figures 17, 18, 19, 20 and 21 in the Appendix for the other insects and the hummingbird.
@@
In our results, $\Omega=2\pi f$ was used for all simulations, and the values for each insect’s flapping frequency are provided in Table 1. Other parameters necessary to implement our natural hovering ES is given in Table 3.
@@ III.2 Re-examining Some Literature Results from the Eyes of Natural Hovering Extremum Seeking
The first topic: re-interpretation of open-loop stability in hovering. It is widely believed in the literature (see Subsection I.1) that hovering of flapping insects (or hummingbirds) is inherently unstable in the open-loop sense due to the pitching motion. However, recently, the authors in [32] argued that open-loop stability can be shown in flapping insects while including the pitching motion for the hawkmoth. They also acknowledged [33] the plausibility of stability being achieved due to some feedback mechanisms similar to what was asserted in many other works in literature (see relevant discussion and citations in Subsection III.C). While we anticipate that natural hovering ES can potentially lead to stable hovering even when pitching motion is included, we focus our attention here on open-loop stability as presented in [69] since we borrowed our model from that work. It is important to note here that there are two kinds of stability that can be achieved in open-loop setting [69] (similar observations also hold even when pitching motion [32] is included). First, open-loop stability for $w=\text{constant}$. This kind of stability does not lead to hovering balance since $w=\text{constant}$ translates into linearly ascending/descending $z$ as noted in multiple figures in [69, 32]. Second, open-loop stability for $w=0$, which leads to hovering balance since $w=0$ translates into constant altitude, i.e., $z=constant$ as expected in stable hovering. As a result, it is quite challenging to achieve precise stable hovering in an open-loop sense because one has to find a trim condition characterizing a flapping amplitude that leads to exact stable hovering condition (i.e., $w=0$), which can be hard to find and requires advanced methods of numerical treatments/approaches as done in [69]. Any slight deviation from said trim condition, will lead to imprecise hovering in the altitude $z$. On the contrary, our natural hovering ES even for the hawkmoth case (see Figure 5), stabilizes about $w=0$ even under disturbances, which always achieves a constant $z$ as naturally observed in hovering. The disparity between some of the results in [69, Figures 3,4,5,6] and our results (see Figure 5) is due to the breaking (de-activation) of the sensation feedback; this leads to a system – even if stabilized in the open-loop sense – that is challenging to maintain stability about the required hovering condition. This is the opposite of the natural hovering ES results. Although this should be obvious, we note that the open-loop approach requires the use of kinematic equation representing flapping; if we set $K=0$ (see Figure 4) in the natural hovering ES, we get an open-loop system as the modulation signal is then the kinematic equation itself that represent flapping. To clarify further our re-interpretation of the open-loop stability, we conducted a simulation comparing between the open-loop case from [69] by using exact $U_{trim}$ value (amplitude of flapping) computed numerically from optimized shooting method and our natural hovering ES with closed-loop ES feedback system (i.e., $K>0$). We used the objective function $J=z^{2}$ for the natural hovering ES (closed-loop). The simulation results (see Figure 7) shows that the trajectory of the ES closed-loop feedback system diverges away from its alignment with the open-loop trajectory, seeking stabilization that achieves hovering (oscillating about $w=0$ and constant $z$). However, the open-loop trajectory is struggling to maintain $z=constant$ even with very small numerical errors in the computation of hovering balance flapping amplitude. With lack of feedback, this slight divergence of $z$ cannot be corrected.
@@ III.3 Comparison Between PID and Natural Hovering Extremum Seeking Approach
In this Subsection, we move forward with a comparison between the proposed natural hovering ES system and Proportional-Integral-Derivative (PID) approach for controlled hovering. The reason we choose a comparison with PID controller is that it was used in hovering literature (eg., [45, 16]), but it is also a traditional simple controller that has the potential of operating in a model-free fashion with low computational cost. We believe this comparison is important as it highlights for many readers, especially those not familiar with control theory/systems, the advantages and robustness of the proposed natural hovering ES system.
PID controllers rely on error feedback from a predefined reference signal and are widely, and most effectively, used for linear time-invariant (LTI) systems where they can be considered model-free and have much larger tolerance in tuning their gains. However, since flapping-wing dynamics are nonlinear, time-varying systems, PID controllers cannot be applied directly or naturally on such systems. Consequently, applying PID control to flapping systems requires multiple approximations and extensions known in control theory and systems. In fact, it is a general knowledge in control theory and systems that PID control application in highly nonlinear systems is suboptimal and lead to gains that are: (1) dependent on a predefined equilibrium and linearization of the system around said equilibrium; and (2) very sensitive to tuning with sometimes very small margin of tolerance. It is worth noting that flapping-wing systems are not just highly nonlinear systems, but they are time-varying as well, which adds another layer of complexity to the problem of controlling hovering via PID-based technique.
@@
We obtained the PID gains as error proportion gain $P=3.66e-03$, integral of error gain $I=2.936e-03$, and derivative of error gain $D=2.00e-03$. We would like to report that the small derivative of error gain $D$ was added to reduce overshoot even though it did not come out naturally from the tuning process. We also would like to report that the system is very sensitive to the gains. Using these gains, the system responded well. The response of the PID-controlled hovering versus the natural hovering ES system is shown in Figure 8, which shows that the natural hovering ES system is performing better in terms of settling time and its lesser overshoot. For the full details of the PID tuning methodology and derivations, we provide the implementation files in [82].
@@
The key advantages of the natural hovering ES system over PID control stem from its model-free, real-time adaptability. Unlike PID when it is applied to nonlinear systems, which requires a linearized transfer function, ES does not rely on explicit system knowledge by any mean. Additionally, as discussed earlier, natural hovering ES does not require an additional input to control the system, but the flapping oscillation itself is the control input. Additionally, PID control struggles with unknown objectives (even if it is able to handle model-free dynamics), as it requires a predefined reference altitude $z_{\text{desired}}$, whereas ES autonomously seeks the optimal hovering condition corresponding to a maximum or minimum of an unknown objective function (refer to earlier discussions about light source seeking). That is, ES is capable of performing real-time optimization and source seeking, unlike PID, which relies purely on error minimization. lastly, PID controllers require precise tuning that may need to be re-adjusted when controlling nonlinear systems, especially if the controller is stabilizing about different equilibrium (altitudes) from a time to another, whereas ES naturally adapts in real-time without requiring manual gain adjustments. To summarize the fundamental differences between PID control and the natural hovering ES system, we present a comparative analysis in Table 5.
@@
Knowledge of the dynamic system model May require an explicit/partial knowledge for the dynamic system model if it is nonlinear. Entirely model-free.
@@
Handling of unknown objectives Requires access to predefined reference signal $z_{\text{desired}}.$ It does not require explicit knowledge of the objective function (including its maximum or minimum information).
@@ III.4 Effect of Delay and Disturbance on the Natural Hovering ES System
In real-world, delays and disturbances/noises are inevitable and can significantly impact the performance of any control system. However, ES systems are usually able to accommodate stochasticity in input or output (feedback measurements), including delays – see for example [83, 84, 85, 86] and some experimental results involving ES systems in noisy real-world robotic applications of light source seeking [78, 79]. In this Subsection, we demonstrate the robustness of the natural hovering ES system under some uncertain conditions, namely the effects of time delay and noise in feedback measurements.
@@
Biological flight systems are constantly exposed to environmental disturbances, including wind gusts and air turbulence, among others. To evaluate the robustness of the natural hovering ES system against such uncertainties and persistent noise condition, we apply noise to the measurement (sensation) of the objective function in a similar fashion to [64, Figure 12]. we introduce a Gaussian noise disturbance added to the feedback altitude $z$, which affects the objective function $J=z^{2}$, with a variance of $\sigma^{2}=0.0001$. The system response under no disturbance and with added noise is illustrated in Figure 9 (bottom). The natural hovering ES system remains stable even under noisy conditions, but small fluctuations are observed in the altitude due to the external perturbations. Despite the persistent noise, the system continues to hover within an acceptable altitude range compared to the case with no noise.
@@ IV Stability analysis of our natural hovering extremum seeking system and comparison with open-loop
Even though ES systems generally have stability guarantees and are known for their robustness [60], it is important to note that our natural hovering ES system needs more investigation to assert its stability. That is, our natural hovering ES system (11) has some difference when compared to the classic class of ES systems in [77, 58]. In particular, in the natural hovering ES system (11), the modulation and demodulation signals have large amplitudes (i.e., scaled by $\Omega$) which is not the case in the classic ES structure – see Figure 4 as opposed to the classic structure [77, Figure 1]. Hence, we proceed with our own stability analysis to the natural hovering ES system (11). The flight dynamics of hovering flapping-wing of insects have been traditionally treated as a nonlinear, time-periodic (NLTP) system [1], where the equilibrium state is represented by a periodic orbit that oscillates about a fixed point equilibrium. Not surprising, averaging techniques [29] have been the standard method in analyzing the NLTP system representing flapping flight and hovering. Averaging techniques are particularly appealing in such analysis because they are an analytical approach, which provides better physical understanding of flapping flight phenomena and the possibility of mimicking control designs. We note that our natural hovering ES system (11) is too, an NLTP system. However, for averaging techniques [29] to apply, the NLTP system (even if it is in a control-affine form [67]) needs to satisfy the averaging canonical form $\dot{\bm{x}}=\epsilon\bm{g}(\bm{x},t)$, where $\epsilon$ is a small parameter, $\bm{g}$ is smooth in $x$ and a $T$-periodic vector field in $t$. Now, does our natural hovering ES system satisfies said averaging canonical form?
@@
while $\tau_{\phi}=\,\Omega\cos(\Omega t)$. It is clear that the system (12) cannot be written in the averaging canonical form, deeming standard averaging techniques [29, 67](e.g., direct/first-order or higher-order averaging) inapplicable. However, the system (12) satisfies the Variation of Constants (VOC) condition in [66, 29]. Clearly, the NLTP system (12) (and hence our natural hovering ES system (11)) can be written as:
@@
by taking $\epsilon=1/\Omega$ and $\bm{Y}\left(\bm{x},\Omega t\right)=\bm{G}(\bm{x})cos(\Omega t)$. As a result of the VOC condition [66, 29] being satisfied, an averaged system to (13) can be derived (see examples of such derivation in relevant applications [69, 74, 73]) and used for stability analysis:
@@
A key limitation in applying VOC-based averaging via (14) is that smothness in $\bm{x}$ is not satisfied in the vector field $\bm{Z}(\bm{x})$ due to the presence of nonsmooth function, namely $|\dot{\phi}|$ (see (12)). This lack of smoothness curtails the use of differentiation to compute the Lie bracket in (14). To address this, we introduce a smooth approximation for the absolute value function $|\dot{\phi}|$ similar to [73]. We approximate the absolute value function as: $|\dot{\phi}|\approx\dot{\phi}h(\dot{\phi})$, where $h(\dot{\phi})=\frac{2}{\pi}\tan^{-1}(n\dot{\phi})$. The larger the $n$, the better the smoothing approximation. Through our analysis, we found that using $n=50$ works well for all insects and hummingbird, offering a sufficiently accurate approximation. Furthermore, Figure 10 shows a precise match between the actual nonsmooth ES dynamics per (11) of the hawkmoth and the smoothed ES dynamics with $n=50$, specifically for the states $w$ and $\dot{\phi}$ starting the simulation with $w_{0}=0$. In said simulation, we used $J=z^{2}$.
@@
After considering the aforementioned smoothing approximation, we are ready to apply the VOC-based averaging in (14) and analyze stability via linearization and eigenvalues, which in turn characterizes the stability of the natural hovering ES system (11). The reader is referred to the code in [82] for the computations involved in (14), linearization about hovering equilibrium, and eigenvalues computation. However, here we present the end result, provided in Table 6. Additionally, we made computations and included also in Table 6 the eigenvalues of stable open-loop hovering [69]; the reader is encouraged to recall relevant comments/discussions about open-loop stability provided in Subsection III.A. The eigenvalues results confirm the improvement in hovering stability by the natural hovering ES system using different objective functions compared to the open-loop case.
Additionally, to visualize the improvement in hovering stability by the natural hovering ES system compared to the open-loop case, we present a planner distribution of the eigenvalues for the hawkmoth case in Figure 11. This analysis corresponds to the eigenvalues listed in Table 6 and provides further insight into how the ES framework enhances hovering stability. The results confirm that via extremum seeking control (ESC), hovering exhibits eigenvalues with more negative real components (i.e., shifts the eigenvalues further into the left-half complex plane), indicating stronger stability properties and more robustness.
Figure 11: Planner distribution of the eigenvalues of the averaged system for the hawkmoth case under different objective functions, demonstrating enhancing in stability with ESC compared with open-loop case for $w$ and $\dot{\phi}$ states.

### v4

@@ III.1 Results and Simulation
In equilibrium hovering, the velocity $w$ has to be 0 as this translates into constant height/altitude $z$ (see the first differential equation in (11)). Therefore, in our simulations we start with an initial disturbance in $w$, that is $w_{0}=0.2$. This disturbance challenges the natural hovering ES system as it needs to overcome temporary deviations from equilibrium and re-establish hovering equilibrium (i.e., $w=0$), demonstrating stability in real-time. The simulations were performed using two objective functions: $J=z^{2}$, where its minimum represents the hovering condition (constant $z$) and $J=\dot{w}^{2}$, where its minimum represents the hovering condition (lift to weight ratio around one). By employing both objective functions in our simulations, the system demonstrated an impressive ability to re-stabilize itself, with the vertical velocity state $w$ oscillating around zero and the altitude remaining constant over time. This behavior reflects stable return to equilibrium hovering after disturbance. The state variables response of the natural hovering ES system (11) are provided in Figure 4 for the hawkmoth, and Figures 11, 12, 13, 14 and 15 in the Appendix for the other insects and the hummingbird. The left column of each figure corresponds to results under the objective function $J=z^{2}$, while the right column of each figure corresponds to results under the objective function $J=\dot{w}^{2}$. The stable hovering conditions of our simulation results (constant $z$ and lift to weight ratio of one) are further clarified in Figure 5 for hawkmoth and Figures 16, 17, 18, 19 and 20 in the Appendix for the other insects and the hummingbird.
@@
Throughout the simulation results, $\Omega=2\pi f$. The learning-rate gain $K$ used for all simulations is given in Table 3, and the values for each insect’s flapping frequency are provided in Table 1. It is important to highlight that we used the same learning-rate gain $K$ for all simulations across all species. This underscores the fact that the natural hovering ES is robust and requires no gain re-tuning when experiencing large variations (e.g., an individual insect/bird that undergoes wing damage to retain its ability to hover stably despite the wing’s altered force-kinematic relationships). Other parameters necessary to implement our natural hovering ES is given in Table 3.
@@ III.2 Re-examining Some Literature Results from the Eyes of Natural Hovering Extremum Seeking
The first topic: re-interpretation of open-loop stability in hovering. It is widely believed in the literature that hovering of flapping insects (or hummingbirds) is inherently unstable in the open-loop sense due to the pitching motion. However, recently, the authors in [19] argued that open-loop stability can be shown in flapping insects while including the pitching motion for the hawkmoth. They also acknowledged [20] the plausibility of stability being achieved due to some feedback mechanisms similar to what was asserted in many other works in literature (see relevant discussion and citations in Subsection III.C). While we anticipate that natural hovering ES can potentially lead to stable hovering even when pitching motion is included, we focus our attention here on open-loop stability as presented in [46] since we borrowed our model from that work. It is important to note here that there are two kinds of stability that can be achieved in open-loop setting [46] (similar observations also hold even when pitching motion [19] is included). First, open-loop stability for $w=\text{constant}$. This kind of stability does not lead to hovering balance since $w=\text{constant}$ translates into linearly ascending/descending $z$ as noted in multiple figures in [46, 19]. Second, open-loop stability for $w=0$, which leads to hovering balance since $w=0$ translates into constant altitude, i.e., $z=constant$ as expected in stable hovering. As a result, it is quite challenging to achieve precise stable hovering in an open-loop sense because one has to find a trim condition characterizing a flapping amplitude that leads to exact stable hovering condition (i.e., $w=0$), which can be hard to find and requires advanced methods of numerical treatments/approaches as done in [46]. Any slight deviation from said trim condition, will lead to imprecise hovering in the altitude $z$. On the contrary, our natural hovering ES even for the hawkmoth case (see Figure 4), stabilizes about $w=0$ even under disturbances, which always achieves a constant $z$ as naturally observed in hovering. The disparity between some of the results in [46, Figures 3,4,5,6] and our results (see Figure 4) is due to the breaking (de-activation) of the sensation feedback; this leads to a system – even if stabilized in the open-loop sense – that is challenging to maintain stability about the required hovering condition. This is the opposite of the natural hovering ES results. Although this should be obvious, we note that the open-loop approach requires the use of kinematic equation representing flapping; if we set $K=0$ (see Figure 3) in the natural hovering ES, we get an open-loop system as the modulation signal is then the kinematic equation itself that represent flapping. To clarify further our re-interpretation of the open-loop stability, we conducted a simulation comparing between the open-loop case from [46] by using exact $U_{trim}$ value (amplitude of flapping) computed numerically from optimized shooting method and our natural hovering ES with closed-loop ES feedback system (i.e., $K>0$). We used the objective function $J=z^{2}$ for the natural hovering ES (closed-loop). The simulation results (see Figure 6) shows that the trajectory of the ES closed-loop feedback system diverges away from its alignment with the open-loop trajectory, seeking stabilization that achieves hovering (oscillating about $w=0$ and constant $z$). However, the open-loop trajectory is struggling to maintain $z=constant$ even with very small numerical errors in the computation of hovering balance flapping amplitude. With lack of feedback, this slight divergence of $z$ cannot be corrected.
@@ III.3 Comparison Between PID and Natural Hovering Extremum Seeking Approach
In this Subsection, we move forward with a comparison between the proposed natural hovering ES system and Proportional-Integral-Derivative (PID) approach for controlled hovering. The reason we choose a comparison with PID controller is that it was used in hovering literature (eg., [28, 16, 17]), but it is also a traditional simple controller that has the potential of operating in a model-free and real-time fashion with low computational cost. We believe this comparison is important as it highlights for many readers, especially those not familiar with control theory/systems, the advantages and robustness of the proposed natural hovering ES system.
PID controllers rely on error feedback from a predefined reference signal and are widely, and most effectively, used for linear time-invariant (LTI) systems where they can be considered model-free and have much larger tolerance in tuning their gains. However, since flapping-wing dynamics are nonlinear, time-varying systems, PID controllers cannot be applied directly or naturally on such systems. Consequently, applying PID control to flapping systems requires multiple approximations and extensions known in control theory and systems. In fact, it is a general knowledge in control theory and systems that PID control application in highly nonlinear systems is suboptimal and leads to gains that are: (1) dependent on a predefined equilibrium and linearization of the system around said equilibrium; and (2) very sensitive to tuning with sometimes very small margin of tolerance. It is worth noting that flapping-wing systems are not just highly nonlinear systems, but they are time-varying as well, which adds another layer of complexity to the problem of controlling hovering via PID-based technique.
@@
We obtained the PID gains as error proportion gain $k_{p}=3.66e-03$, integral of error gain $k_{i}=2.936e-03$, and derivative of error gain $k_{d}=2.00e-03$. We would like to report that the small derivative of error gain $k_{d}$ was added to reduce overshoot even though it did not come out naturally from the tuning process. We also would like to report that the system is very sensitive to the gains. To illustrate this sensitivity, the initial disturbance $w_{0}$ was varied while the PID gains (mentioned earlier) kept constant. Table 5 compares overshoot and settling time for the resulting PID response against the natural hovering ES controller, which uses the same learning-rate gain $K$ in every case. Even small changes in $w_{0}$ drive the PID overshoot above $7\!\times\!10^{2}\,\%$ and lengthen the settling time to $\approx 10$ s, whereas ES maintains overshoot below $26\%$ and settles in time $\leq 1.6$ s without re-tuning. It should be noted that the settling time and overshoot values were calculated using MATLAB’s built-in (stepinfo) function. To visualize this effect, Figure 7 shows the response of the PID-controlled hovering versus the natural hovering ES system for $w_{0}=0.2$, which shows that the natural hovering ES system is performing better in terms of settling time and its lesser overshoot. These results confirm that the PID controller must be re-optimized for each disturbance magnitude, while the ES law is robust to large variations/disturbances. For the full details of the PID tuning methodology and derivations, we provide the implementation files in [60].
Table 5: Overshoot and settling‐time performance of the ES and PID controllers for five initial perturbations $w_{0}$.
$\bm{w_{0}}$ Overshoot [%] Settling time (s)
ES PID ES PID
0.2 23.63 770.22 1.39 10.22
0.3 25.22 769.85 1.39 10.20
-0.2 22.82 772.45 1.54 10.26
-0.3 23.67 773.28 1.60 10.27
-0.1 22.41 771.70 1.29 10.24
For more demonstrative, stronger, and fairer comparison, we tried to get the optimal PID gains by using the genetic algorithm (GA) optimizer. Running GA for each insect, the optimal PID gains are obtained as shown in Table 6; the GA code and configuration files are available in [60]. To demonstrate the sensitivity of even optimal PID gains to changes in operating conditions, the GA was rerun for the hawkmoth under five different initial $w_{0}$. Table 7 shows that every change in $w_{0}$ yields different $(k_{p},k_{i},k_{d})$ values, confirming that the PID controller must be re-optimized (re-tuned) whenever the operating condition changes.
Table 6: Optimal PID gains obtained by GA for $w_{0}=0.2$
Insect $k_{p}$ $k_{i}$ $k_{d}$
Hawkmoth $8.912\text{e}{-02}$ $2.585\text{e}{-01}$ $2.405\text{e}{-02}$
Bumblebee $5.000\text{e}{-04}$ $4.000\text{e}{-04}$ $1.000\text{e}{-04}$
Cranefly $1.421\text{e}{-04}$ $3.983\text{e}{-04}$ $4.679\text{e}{-05}$
Dragonfly $5.000\text{e}{-04}$ $4.000\text{e}{-04}$ $9.999\text{e}{-05}$
Hoverfly $3.337\text{e}{-04}$ $4.000\text{e}{-04}$ $9.999\text{e}{-05}$
Hummingbird $5.000\text{e}{-02}$ $4.000\text{e}{-02}$ $1.000\text{e}{-02}$
Table 7: Hawkmoth—optimal PID gains identified for five initial disturbances $w_{0}$.
$\bm{w_{0}}$ $\bm{k_{p}}$ $\bm{k_{i}}$ $\bm{k_{d}}$
$0.20$ $8.912\mathrm{e}{-02}$ $2.585\mathrm{e}{-01}$ $2.405\mathrm{e}{-02}$
$0.30$ $1.147\mathrm{e}{-01}$ $3.178\mathrm{e}{-01}$ $2.429\mathrm{e}{-02}$
$-0.10$ $2.411\mathrm{e}{-02}$ $1.027\mathrm{e}{-01}$ $2.296\mathrm{e}{-02}$
$-0.20$ $2.386\mathrm{e}{-02}$ $1.152\mathrm{e}{-01}$ $6.272\mathrm{e}{-03}$
$-0.30$ $2.188\mathrm{e}{-02}$ $8.514\mathrm{e}{-02}$ $4.248\mathrm{e}{-03}$
Moreover, Table 8 compares overshoot and settling time for the natural hovering ES (same gain $K$ in all cases) and the GA–tuned (i.e., optimal) PID controller. The optimal PID improves on the hand-tuned version when compared to its performance in Table 5, yet still requires more time to settle with high overshoot. In contrast, the natural hovering ES remains more efficient and robust with less overshoot and settling time without any re-tuning. To visualize this effect, Figure 8 plots altitude $z(t)$ and velocity $w(t)$ for two representative disturbances, $w_{0}=0.2$ and $w_{0}=-0.2$. Even with their GA-optimized gains (shown in Table(8)), the optimal PID response takes longer to converge and overshoots more than the natural hovering ES, which further confirms the same conclusion about the ES effectiveness and robustness. These results highlight a core advantage of the natural hovering ES controller: a single gain setting delivers consistent performance across a wide range of initial conditions/disturbances, whereas PID requires case-by-case re-optimization with sub-optimal performance/robustness. We provide the implementation files in [60]
Figure 8: Comparison of altitude stabilization using the optimal PID controller (red) versus the natural hovering ES system (black). The ES loop settles faster and with smaller overshoot for both initial conditions ($w_{0}=0.2$ and $w_{0}=-0.2$), while using the same ESC feedback gain in every case and different optimal PID gains as shown in table 7 .
Table 8: Overshoot and settling‐time performance of the ES and optimal PID controllers for five initial perturbations $w_{0}$.
$\bm{w_{0}}$ Overshoot [%] Settling time (s)
ES Optimal PID ES Optimal PID
0.2 23.63 34.94 1.39 2.27
0.3 25.22 29.98 1.39 1.97
-0.2 22.82 142.88 1.54 3.18
-0.3 23.67 178.51 1.60 2.45
-0.1 22.41 68.36 1.29 8.85
@@
Overall, the key advantages of the natural hovering ES system over PID control stem from its model-free, real-time adaptability. Unlike PID when it is applied to nonlinear systems, which requires a linearized transfer function, ES does not rely on explicit system knowledge by any mean. Additionally, as discussed earlier, natural hovering ES does not require an additional input to control the system, but the flapping oscillation itself is the control input. Additionally, PID control struggles with unknown objectives (even if it is able to handle model-free dynamics), as it requires a predefined reference altitude $z_{\text{desired}}$, whereas ES autonomously seeks the optimal hovering condition corresponding to a maximum or minimum of an unknown objective function (refer to earlier discussions about light source seeking). That is, ES is capable of performing real-time optimization and source seeking, unlike PID, which relies purely on error minimization. lastly, PID controllers require precise tuning that may need to be re-adjusted when controlling nonlinear systems, especially if the controller is stabilizing about different equilibrium (altitudes) from a time to another, whereas ES naturally adapts in real-time without requiring manual gain adjustments. To summarize the fundamental differences between PID control and the natural hovering ES system, we present a comparative analysis in Table 9.
@@
Knowledge of the dynamic system model May require an explicit/partial knowledge for the dynamic system model if it is nonlinear, at least for re-tuning. Entirely model-free with fixed leaning-rate gain.
@@
Handling of unknown objectives Requires access to predefined reference signal $z_{\text{desired}}.$. That is, it needs global information. It does not require explicit knowledge of the objective function (including its maximum or minimum information). It operates with local measurements
@@ III.4 Effect of Delay and Disturbance on the Natural Hovering ES System
In real-world, delays and disturbances/noises are inevitable and can significantly impact the performance of any control system. However, ES systems are usually able to accommodate stochasticity in input or output (feedback measurements), including delays – see, for example, some recent developments [61, 62, 63, 64] and some experimental results involving ES systems in noisy real-world robotic applications of light source seeking [56, 57]. In this Subsection, we demonstrate the robustness of the natural hovering ES system under some uncertain conditions, namely the effects of time delay and noise in feedback measurements.
@@
Biological flight systems are constantly exposed to environmental disturbances, including wind gusts and air turbulence, among others. To evaluate the robustness of the natural hovering ES system against such uncertainties and persistent noise condition, we apply noise to the measurement (sensation) of the objective function in a similar fashion to [40, Figure 12]. we introduce a Gaussian noise disturbance added to the feedback altitude $z$, which affects the objective function $J=z^{2}$, with a variance of $\sigma^{2}=0.0001$. The system response under no disturbance and with added noise is illustrated in Figure 9 (bottom). The natural hovering ES system remains stable even under noisy conditions, but small fluctuations are observed in the altitude due to the external perturbations. Despite the persistent noise, the system continues to hover within an acceptable altitude range compared to the case with no noise.
@@ IV Stability analysis of our natural hovering extremum seeking system and comparison with open-loop
Even though ES systems generally have stability guarantees and are known for their robustness [36], it is important to note that our natural hovering ES system needs more investigation to assert its stability. That is, our natural hovering ES system (11) has some difference when compared to the classic class of ES systems in [55, 34]. In particular, in the natural hovering ES system (11), the modulation and demodulation signals have large amplitudes (i.e., scaled by $\Omega$) which is not the case in the classic ES structure – see Figure 3 as opposed to the classic structure [55, Figure 1]. Hence, we proceed with our stability analysis to the natural hovering ES system (11) based on the framework provided in [44]. The flight dynamics of hovering flapping-wing of insects have been traditionally treated as a nonlinear, time-periodic (NLTP) system [1], where the equilibrium state is represented by a periodic orbit that oscillates about a fixed point equilibrium. Not surprising, averaging techniques [66] have been the standard method in analyzing the NLTP system representing flapping flight and hovering. Averaging techniques are particularly appealing in such analysis because they are an analytical approach, which provides better physical understanding of flapping flight phenomena and the possibility of mimicking control designs. We note that our natural hovering ES system (11) is too, an NLTP system. However, for averaging techniques [66] to apply, the NLTP system (even if it is in a control-affine form [43]) needs to satisfy the averaging canonical form $\dot{\bm{x}}=\epsilon\bm{g}(\bm{x},t)$, where $\epsilon$ is a small parameter, $\bm{g}$ is smooth in $x$ and a $T$-periodic vector field in $t$. Now, does our natural hovering ES system satisfies said averaging canonical form?
@@
while $\tau_{\phi}=\,\Omega\cos(\Omega t)$. It is clear that the system (12) cannot be written in the averaging canonical form, deeming standard averaging techniques [66, 43](e.g., direct/first-order or higher-order averaging) inapplicable. However, the system (12) satisfies the Variation of Constants (VOC) condition in [42, 66, 44]. Clearly, the NLTP system (12) (and hence our natural hovering ES system (11)) can be written as:
@@
by taking $\epsilon=1/\Omega$ and $\bm{Y}\left(\bm{x},\Omega t\right)=\bm{G}(\bm{x})cos(\Omega t)$. As a result of the VOC condition [42, 66, 44] being satisfied, an averaged system to (13) can be derived and used for stability analysis:
@@
A key limitation in applying VOC-based averaging via (14) is that smothness in $\bm{x}$ is not satisfied in the vector field $\bm{Z}(\bm{x})$ due to the presence of nonsmooth function, namely $|\dot{\phi}|$ (see (12)). This lack of smoothness curtails the use of differentiation to compute the Lie bracket in (14). To address this, we introduce a smooth approximation for the absolute value function $|\dot{\phi}|$ similar to [50].
We approximate the absolute value function as: $|\dot{\phi}|\approx\dot{\phi}h(\dot{\phi})$, where $h(\dot{\phi})=\frac{2}{\pi}\tan^{-1}(n\dot{\phi})$. The larger the $n$, the better the smoothing approximation. Through our analysis, we found that using $n=50$ works well for all insects and hummingbird, offering a sufficiently accurate approximation. Furthermore, Figure 10 shows a precise match between the actual nonsmooth ES dynamics per (11) of the hawkmoth and the smoothed ES dynamics with $n=50$, specifically for the states $w$ and $\dot{\phi}$ starting the simulation with $w_{0}=0$. In said simulation, we used $J=z^{2}$.